# auxmulticols[1]

### v1

@@ Abstract.
journalyear: 2023journalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXXDisclaimer: The information and views set out in this article are those of the authors and do not necessarily reflect the official opinion of the institutions.
@@ 1. Introduction
These regimes, now a regulatory standard enacted in numerous global privacy laws (201), primarily follow the European Union General Data Protection Regulation (GDPR) model, which grants the most extensive set of data rights (201; 91) (see Appendix Tab. 1). The U.S. has adopted similar approaches, as seen in the California Consumer Privacy Protection Act (CCPA), modeled after the GDPR.
The proposition that bestowing individuals with data rights can mitigate corporate malfeasance and safeguard personal privacy is a subject of disagreement (201; 233; 123). Many hold a positive position on rights regimes, but the traditional belief hinges on an intricate, contingent network of assumptions regarding user awareness, the capacity to exercise their rights, and the companies’ diligence in offering, elucidating, and applying controls.
@@ 2. Data Rights Assessment Framework & Method
Since the European data rights model has become the standard in the field (201; 233; 91), we select the GDPR data rights in Tab. 1 as the focus of our analysis. We developed a general data rights assessment framework consisting of 15 key questions (see Appendix A) by synthesizing 201 interdisciplinary empirical materials. This framework provides a comprehensive evaluation of rights-based privacy approaches, considering the perspectives of the primary actors in the information economy: individuals, companies or developers, and regulators. Practitioners and researchers can use our three pronged framework as a starting point to analyze other data-rights frameworks like CCPA.
Based on the extant empirical studies, we discuss the 15 questions for the three actors (users, companies, and regulators) and reveal open challenges in §3, §4, and §5. For each question, we examine whether there is divergent evidence and, if so, why. Tab. 3 summarizes the evaluation results.
@@ 3. Users’ perspective
(U1.1) Disparity in data rights knowledge. The understanding of data rights is not uniform, with some rights being more recognized than others (67; 184; 119). A large-scale user survey by Eurobarometer (67) reveals that the least known rights are the right to data portability and the right to avoid automated decision making. However, the rights to access, erasure, and object are relatively well-known. Consistent results emerge from a German study (119).
@@
(U1.3) Contextual variation in knowledge. The awareness level also changes dramatically depending on the technology context (122; 181; 128; 210). For instance, many users understand how to exercise the right to access with smart speakers like Amazon Echo and Google Home (122), yet only 24% of patients are aware of accessing their health information (181). Moreover, many users are uninformed about the availability of privacy policies of their voice apps (128). Recall the fact that the right to information and access are well-known in general (164; 67; 184; 119). As a result, this discrepancy highlights the gap between abstract knowledge and practical application of data rights.
@@
(U2.1) Positive perception. Across numerous contexts, we find that users often hold a positive perception and express interest in their data rights, ranging from the right to information to the right to avoid automated decision making (202; 174; 122; 181; 159; 38; 7; 175; 113; 86; 120; 107; 66; 245; 244). For instance, users appreciate being informed about their data rights by service providers (38; 245; 244). They often express curiosity about the reasons for data collection and how this data is utilized by firms (122; 86; 120; 152; 34; 138; 60). Indeed, depending on the context, users want different types of information and different notification frequencies (63; 244). In an interview regarding a retail loyalty card program, some consumers considered it beneficial to access the personal information collected about them (7). Similarly, patients acknowledged the value of the right to access and expressed a desire to exercise this right (181). Hence, many users perceive the right to information and access positively and wish to exercise these rights.
@@
The desire to exercise the right to erasure is also prevalent (7; 152; 34; 211). For example, despite being under a legal surveillance regime, some surveilled subjects (or their families) felt they should have the right to delete their data (159). Approximately 89% of users in another study wanted the ability to delete their data and verify the erasure (135). A favorable view is also held for the right to object. Some survey participants appreciated the ability to opt-out of receiving direct marketing (175). Moreover, users acknowledged the importance of the right to avoid automated decision-making (174; 175; 113; 192; 221). For example, some users preferred making independent choices, even if they were merely suggestions or recommendations on a website (192). The desire to avoid automated decision-making heightened when it involved personal matters like insurance, mortgage loans, or exam results evaluation (174; 175; 113; 221).
@@
(U2.2) Negative perception. However, not all users desire data rights. Some individuals do not feel the need to control their data and show no interest in data rights (181; 7; 113; 108; 62; 187; 152). In fact, a survey shows that, on average, users have a neutral willingness to exercise data rights (203). This is counterintuitive since people generally value their privacy (116; 2). We explore the reasons behind this lack of desire for data rights:
@@
(i) Trust: Trust in companies holding personal data impacts users’ perceptions of data rights (108; 187). Users who have high trust in a company are less likely to feel the need to read privacy policies or to be informed about the company’s data practices (108; 187).
@@
(ii) Indifference: Some users exhibit indifference or insensitivity towards the use of their data, which diminishes their interest in data rights (181; 175; 113; 66). As an example, a small fraction of patients in a study (about 5%) did not find it problematic for others to access their health information at any time, thus showing no interest in their data rights (181). Other individuals perceive that personalized ads and direct marketing do not harm their privacy significantly; hence, they do not aspire to exercise their data rights (175): “I will most likely not pursue this [the right to erasure]. From the advertisements that Facebook continuously shows me, I draw the conclusion that they do not really know too much about me as a person.” Further, a minority of participants in a user experiment (2 out of 16) were not concerned about the transparency of a systems (66).
@@
(iii) Privacy fatigue: Another key reason for users’ lack of desire for data rights is the perceived burden and cost of exercising these rights (7; 62; 188; 69; 229). Some users deem the explanations of automated decision-making boring and irrelevant (113). The constant need to exercise data rights is particularly undesirable, with very few expressing a wish for such continuous engagement (7; 245). Even though some individuals recognize the importance of data rights, they find that active engagement leads to privacy fatigue and annoyance (62; 188; 229). This burden leads some users to prefer delegating their data rights to others, such as automatic tools or third parties, rather than managing their data directly (229; 138). Interestingly, even with accessible online tools that simplify the exercise of data rights, some users still perceive the process as too time-consuming (69; 228): “I have better things to do with my time, frankly, than to be reviewing this.”
@@
(iv) Fear and low understanding: Additionally, some users may not desire data rights due to heightened fear from learning more about personal data processing, or due to a lack of understanding about the benefits of data rights (7; 176). For instance, the right to data portability is often undervalued due to a lack of understanding (175; 176; 189).
@@
(U3.1) Attitude vs. actual behavior Even if users have a significant desire for data rights, they rarely exercise them in practice (175; 120; 176; 219; 142). For instance, a survey indicates that while many express the desire to understand how companies utilize their cookies, few actually read the privacy policy to obtain this information (120). A striking disparity was observed with GDPR implementation; many users who previously expressed strong intentions to exercise data rights did not follow through (176). Specifically, for the right to delete, about 40% of the users claimed that they will “definitely” use this right before enforcing the GDPR, but the following survey conducted after 1 and 2 years shows that only about 15% had exercised the right (176). This incongruity is consistent with the well-known “privacy paradox”—a discrepancy between privacy concerns and actual behavior (116; 2).
(U3.2) Users’ non-exercise of data rights. The reality is that many users do not exercise GDPR data rights. It is common for people to quickly scroll through without actually reading privacy policies (128; 175; 142; 16; 22; 215; 154; 166). User experiments on a tracking app and a social networking service found that over 90% of participants spent less than 30 seconds on the privacy policy page (22; 154), and it seems that only 1% of people read the entire text of privacy policies (175).
Moreover, the lack of exercising rights is not limited to the right to information; it extends to other rights, including the right to erasure. Most Europeans (70% of survey respondents (135)) have never applied for data deletion, and only 1.13% frequently delete their cookie data (161). Service providers also report a lack of requests for data access and erasure (7; 219; 166). Some companies claim there had been no user access requests and less than ten profile deletion requests over a full year, contrary to their expectations (219; 166). Additionally, data portability is infrequently utilized; only 7% of users surveyed have exercised this right (175). Therefore, it is not common for users to exercise GDPR data rights.
@@
In contexts involving their children (e.g., smart homes), users can also pay more attention to privacy policies in smart devices surveilling their homes (211); certain parents consider privacy policies crucial to their children’s safety when selecting smart devices (211). Furthermore, some users regularly examine and delete personal data collected by smart devices related to their daily lives (122; 142; 41).
@@
Data rights have a significant influence on users’ perceived privacy concerns and uncertainties. Specifically, providing users with higher, simpler control over their data can mitigate their perceived privacy concerns, risks, and uncertainties (237; 90; 69; 120). A user experiment, for instance, demonstrates that showing users the personal information they have shared can decrease perceived privacy risks compared to merely viewing a privacy policy (90). When the system further allowed users to delete and modify their data, perceived privacy concerns and risks decreased.
@@
However, this observation contradicts previous findings (35; 146; 42) suggesting that providing users with data control can paradoxically lead to more data disclosure, a phenomenon known as the Peltzman effect where people tend to take riskier actions under security measures. While the exact reason for this contradiction is not clear, one hypothesis is that the mention of GDPR itself stimulated privacy-seeking behavior among users in the study (160). The only major difference across these studies is the mention of GDPR, and it’s possible that the priming effect of the term GDPR outweighed the Peltzman effect. Relatedly, one study (42) reports that granting data control to users did not significantly alter their data-sharing behavior; most participants maintained their willingness not to share personal information even after being given data control. Furthermore, evidence shows that users are more inclined to disclose their health data when given a reason for data collection that highlights the benefits of the product (21). In conclusion, further research is necessary to understand the impact of GDPR data rights on data disclosure behavior.
@@ 4. Companies’ perspective
A significant number of companies demonstrate a limited understanding of data rights (243; 149; 151). A user survey of businesses across eight EU countries revealed that only approximately 50-65% of participants correctly answered questions about the right to access, the right to erasure, and the right to object (243). Additionally, some app developers mistakenly believe that the GDPR is inapplicable to them because their primary market is outside the EU (149). Similarly, several startups misinterpret the GDPR, asserting that certain data rights cannot apply to their key technologies (151). Contrary to their beliefs, the GDPR applies to all businesses collecting personal data from citizens of European Economic Area (EEA) countries, irrespective of their primary market and business sectors. Notably, studies by Zanker et al. (243) and Nguyen et al. (149) conducted in early 2020 and early 2021, respectively, highlight the persistent deficiency in companies’ understanding of GDPR data rights even 2-3 years after its implementation.
@@
(C2.2) Willingness to implement data rights. Not all developers disregard users’ data rights. Many service providers emphasize the importance of enforcing them (95; 247; 8; 74; 134; 133; 81). They highlight the growing user awareness of their rights and believe the exercise of data rights will gain prominence with the adoption of GDPR (95; 247). Most European business leaders (about 86%) recognize the importance of GDPR compliance (74), a figure significantly higher than average individuals fully understanding GDPR data rights (50-65%) (243). This indicates that service providers can appreciate the importance of GDPR, even with a partial understanding of the law. Reasons for compliance include fear of fines and respect for users’ data rights (74). Likewise, practitioners knowledgeable in ML showed appreciation for the right to get a human review in automated decision-making and stated the importance of ML’s transparency (134; 133).
@@
Adherence to the GDPR data rights necessitates significant effort from companies, prompting various internal measures such as employee training, consultation procurement, and data management process modifications (243; 74; 92; 95). A survey reveals that approximately 91% of European companies have conducted at least one employee training session on this subject (243). Numerous companies report engaging in GDPR compliance consultation services (74). In preparation for the GDPR, most companies have incurred costs between €1,000 and €50,000, predominantly for employee education and consultancy services (243; 74).
@@
Several companies have implemented substantial changes to their data management processes to comply with data rights (92; 95; 105). An interview with a publicly listed ML company revealed significant changes to their data handling practices over six months, guided by their legal team to ensure GDPR compliance (92). Enterprise Architecture Management (EAM), which provides companies with strategies for efficient development, has played a crucial role in implementing data rights (95). For instance, some EAMs have defined procedures for developers to follow when implementing data rights and have specified the correct communication channels for notification.
@@
(C4.1) Right to information. A plethora of studies have scrutinized the compliance of privacy policies with the GDPR across various platforms such as web, mobile applications, and IoT (4; 88; 129; 242; 190; 225; 157; 208; 37; 68; 143; 219; 222; 145; 163; 212; 191; 9; 128; 246; 197; 213; 214; 155; 24; 5; 57; 23; 32; 93; 141; 114). These studies employ manual analyses or automated tools, such as Machine Learning (ML) or Natural Language Processing (NLP), to scrutinize privacy policies. Although methodologies and evaluation criteria vary across studies, the consensus is that few privacy policies currently meet all GDPR requirements.
@@
A noticeable disparity exists in the compliance rates for individual GDPR requirements. Over 90% of policies outline the collection of personal data and its purposes, though some instances of incomplete information have been reported (129; 242; 68; 246; 114). However, the majority of policies do not describe data rights like access and erasure (129; 32) even though European websites tend to state more extensive data rights than websites from other countries (240). Furthermore, many policies struggle to accurately inform users about the flow and transfer of data (222; 246; 209; 110; 141). A significant difference in compliance rates across websites in various countries is also found; European websites tend to have a higher compliance rate than others (240; 93).
@@
Surprisingly, some services do not provide a privacy policy at all (128; 197; 222; 156; 155; 141; 140). The availability of privacy policies varies significantly across platforms. For instance, most websites (4,566 out of 5,045 surveyed) provide a privacy policy (156). Yet, a mere 16% of 6,843 analyzed pornographic websites have a privacy policy (222). Similarly, only 62% of Google Assistant actions, which serve as voice-assistance applications, provide a privacy policy (128). Moreover, more than half of Greek and UK and more than one-third of German websites did not include a cookie notice (110; 117), and 12 out of 30 fertility apps analyzed did not prompt privacy notices (141).
@@
(C4.2) Rights to access, data portability, erasure, and object. A multitude of studies have delved into the extent to which systems uphold the right to access and data portability, typically by issuing a direct request (218; 226; 216; 39; 118; 204; 219; 220; 239; 248). These studies generally reveal that the majority of requests are honored within the one-month GDPR deadline. Similarly, an interview indicated that over 90% of service providers processed all requests for access and erasure within a month (166). However, compliance rates can be low in specific sectors (118; 220); for instance, a 2019 study showed that only 41% of popular German mobile apps provided a copy of personal data within a month (220).
@@
Moreover, past studies frequently observe that the data procured through the access right are incomplete (226; 216; 220; 118; 169); the data may not include certain personal information that users supplied. Furthermore, while many systems allow users to download their personal data in a machine-readable format, few provide the data import features required for data portability (119; 218). There are examples of systems rejecting users’ requests to download their data without providing a justification (248).
@@
Numerous technical and managerial challenges can make it virtually impossible for businesses to implement data rights. For instance, some suggest that, for efficiency, companies should deploy a system managing all applications processing personal information, as users can make daily requests to control their data (95). However, this would be challenging for applications run by independent departments (95). Furthermore, budget constraints can impede the implementation of data rights, as companies need to invest significantly in employee training and IT solutions for data management (61; 199).
(C5.1) Right to erasure. Among the eight GDPR rights, the right to erasure was most frequently cited as challenging by developers (151; 11; 135; 31; 92; 95; 44; 171). Removing user data, especially from backups and archives, appears to be a complex, costly, and unclear process (95; 44; 135). When backups need to be restored, previously deleted data may resurface, suggesting a failure in the erasure process. Moreover, larger enterprises tend to struggle more with this right due to the sheer volume of data (171). Some even admit that applying this right in their systems is impossible (11).
Confusion amplifies when considering ML (31; 31; 92). There is a lack of clear guidelines on whether to remove deleted data from all model training, test, and validation sets, and whether to delete the model itself upon receiving an erasure request from users. Another aspect is a conflict between data rights and technologies’ goals. Representatively, blockchain companies suffer from this problem (151). Given the immutability of blockchain technology, erasing data stored in the system is inherently impossible. Many blockchain companies discuss this issue in their privacy policies, with some stating outright that exercising this right is impossible (186). However, a few claims they can delete personal information from their blockchain systems, but details on how this is achieved are typically undisclosed (186).
@@
(C5.2) Right to information and access. Companies often find the rights to information and access challenging to implement (149; 11; 219). Firms must provide accurate information about data processing to users, necessitating a comprehensive understanding of data flows and processing mechanisms. However, many companies struggle with this, especially when third parties involved in online advertising also process the data. Full awareness of how these third parties handle user data can be a formidable task. Several studies reveal that most developers grapple with this lack of knowledge (149; 219; 125).
Moreover, the right to information poses a challenge in ML systems (92; 77; 58; 30). The right to information necessitates that corporations disclose the underlying logic of algorithms that may have consequential or legal effects on users. This disclosure, however, may not be feasible due to the inherent “black box” nature of ML algorithms. There are extant techniques, such as LIME (183), devised to demystify this issue, yet the resultant explanations are often technically complex and inaccessible to users lacking comprehensive ML knowledge (58; 30). This complexity often leads engineers to use these interpretability methods primarily to enhance ML system efficiency, deviating from the original intent of explicating algorithms to users (30).
Moreover, tensions may emerge between data rights and corporate interests, as ML algorithms can be considered proprietary to a company. This poses a critical question: To what extent should companies be obligated to reveal their algorithms (58)? A high degree of ML transparency could paradoxically infringe on user privacy by potentially revealing user identities embedded within the training dataset(58; 30).

### v2

@@ Abstract.
journalyear: 2025journalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXXDisclaimer: The information and views set out in this article are those of the authors and do not necessarily reflect the official opinion of the institutions.
@@ 1. Introduction
These regimes, now a regulatory standard enacted in numerous global privacy laws (213), primarily follow the European Union General Data Protection Regulation (GDPR) model, which grants the most extensive set of data rights (213; 96) (see Appendix Tab. 1). The U.S. has adopted similar approaches, as seen in the California Consumer Privacy Protection Act (CCPA), modeled after the GDPR.
The proposition that bestowing individuals with data rights can mitigate corporate malfeasance and safeguard personal privacy is a subject of disagreement (213; 245; 131). Many hold a positive position on rights regimes, but the traditional belief hinges on an intricate, contingent network of assumptions regarding user awareness, the capacity to exercise their rights, and the companies’ diligence in offering, elucidating, and applying controls.
@@ 2. Data Rights Assessment Framework & Method
Since the European data rights model has become the standard in the field (213; 245; 96), we select the GDPR data rights in Tab. 1 as the focus of our analysis. We developed a general data rights assessment framework consisting of 15 key questions (see Appendix A) by synthesizing 201 interdisciplinary empirical materials. This framework provides a comprehensive evaluation of rights-based privacy approaches, considering the perspectives of the primary actors in the information economy: individuals, companies or developers, and regulators. Practitioners and researchers can use our three pronged framework as a starting point to analyze other data-rights frameworks like CCPA.
Based on the extant empirical studies, we discuss the 15 questions for the three actors (users, companies, and regulators) and reveal open challenges in §3, §4, and §5. For each question, we examine whether there is divergent evidence and, if so, why. Tab. LABEL:tab:framework summarizes the evaluation results.
@@ 3. Users’ perspective
(U1.1) Disparity in data rights knowledge. The understanding of data rights is not uniform, with some rights being more recognized than others (72; 195; 127). A large-scale user survey by Eurobarometer (72) reveals that the least known rights are the right to data portability and the right to avoid automated decision making. However, the rights to access, erasure, and object are relatively well-known. Consistent results emerge from a German study (127).
@@
(U1.3) Contextual variation in knowledge. The awareness level also changes dramatically depending on the technology context (130; 192; 136; 222). For instance, many users understand how to exercise the right to access with smart speakers like Amazon Echo and Google Home (130), yet only 24% of patients are aware of accessing their health information (192). Moreover, many users are uninformed about the availability of privacy policies of their voice apps (136). Recall the fact that the right to information and access are well-known in general (175; 72; 195; 127). As a result, this discrepancy highlights the gap between abstract knowledge and practical application of data rights.
@@
(U2.1) Positive perception. Across numerous contexts, we find that users often hold a positive perception and express interest in their data rights, ranging from the right to information to the right to avoid automated decision making (214; 185; 130; 192; 170; 41; 8; 186; 121; 91; 128; 114; 71; 259; 258). For instance, users appreciate being informed about their data rights by service providers (41; 259; 258). They often express curiosity about the reasons for data collection and how this data is utilized by firms (130; 91; 128; 162; 37; 147; 64). Indeed, depending on the context, users want different types of information and different notification frequencies (68; 258). In an interview regarding a retail loyalty card program, some consumers considered it beneficial to access the personal information collected about them (8). Similarly, patients acknowledged the value of the right to access and expressed a desire to exercise this right (192). Hence, many users perceive the right to information and access positively and wish to exercise these rights.
@@
The desire to exercise the right to erasure is also prevalent (8; 162; 37; 223). For example, despite being under a legal surveillance regime, some surveilled subjects (or their families) felt they should have the right to delete their data (170). Approximately 89% of users in another study wanted the ability to delete their data and verify the erasure (144). A favorable view is also held for the right to object. Some survey participants appreciated the ability to opt-out of receiving direct marketing (186). Moreover, users acknowledged the importance of the right to avoid automated decision-making (185; 186; 121; 203; 233). For example, some users preferred making independent choices, even if they were merely suggestions or recommendations on a website (203). The desire to avoid automated decision-making heightened when it involved personal matters like insurance, mortgage loans, or exam results evaluation (185; 186; 121; 233).
@@
(U2.2) Negative perception. However, not all users desire data rights. Some individuals do not feel the need to control their data and show no interest in data rights (192; 8; 121; 115; 66; 198; 162). In fact, a survey shows that, on average, users have a neutral willingness to exercise data rights (215). This is counterintuitive since people generally value their privacy (124; 2). We explore the reasons behind this lack of desire for data rights:
@@
(i) Trust: Trust in companies holding personal data impacts users’ perceptions of data rights (115; 198). Users who have high trust in a company are less likely to feel the need to read privacy policies or to be informed about the company’s data practices (115; 198).
@@
(ii) Indifference: Some users exhibit indifference or insensitivity towards the use of their data, which diminishes their interest in data rights (192; 186; 121; 71). As an example, a small fraction of patients in a study (about 5%) did not find it problematic for others to access their health information at any time, thus showing no interest in their data rights (192). Other individuals perceive that personalized ads and direct marketing do not harm their privacy significantly; hence, they do not aspire to exercise their data rights (186): “I will most likely not pursue this [the right to erasure]. From the advertisements that Facebook continuously shows me, I draw the conclusion that they do not really know too much about me as a person.” Further, a minority of participants in a user experiment (2 out of 16) were not concerned about the transparency of a systems (71).
@@
(iii) Privacy fatigue: Another key reason for users’ lack of desire for data rights is the perceived burden and cost of exercising these rights (8; 66; 199; 74; 241). Some users deem the explanations of automated decision-making boring and irrelevant (121). The constant need to exercise data rights is particularly undesirable, with very few expressing a wish for such continuous engagement (8; 259). Even though some individuals recognize the importance of data rights, they find that active engagement leads to privacy fatigue and annoyance (66; 199; 241). This burden leads some users to prefer delegating their data rights to others, such as automatic tools or third parties, rather than managing their data directly (241; 147). Interestingly, even with accessible online tools that simplify the exercise of data rights, some users still perceive the process as too time-consuming (74; 240): “I have better things to do with my time, frankly, than to be reviewing this.”
@@
(iv) Fear and low understanding: Additionally, some users may not desire data rights due to heightened fear from learning more about personal data processing, or due to a lack of understanding about the benefits of data rights (8; 187). For instance, the right to data portability is often undervalued due to a lack of understanding (186; 187; 200).
@@
(U3.1) Attitude vs. actual behavior Even if users have a significant desire for data rights, they rarely exercise them in practice (186; 128; 187; 231; 151). For instance, a survey indicates that while many express the desire to understand how companies utilize their cookies, few actually read the privacy policy to obtain this information (128). A striking disparity was observed with GDPR implementation; many users who previously expressed strong intentions to exercise data rights did not follow through (187). Specifically, for the right to delete, about 40% of the users claimed that they will “definitely” use this right before enforcing the GDPR, but the following survey conducted after 1 and 2 years shows that only about 15% had exercised the right (187). This incongruity is consistent with the well-known “privacy paradox”—a discrepancy between privacy concerns and actual behavior (124; 2).
(U3.2) Users’ non-exercise of data rights. The reality is that many users do not exercise GDPR data rights. It is common for people to quickly scroll through without actually reading privacy policies (136; 186; 151; 19; 25; 227; 164; 177). User experiments on a tracking app and a social networking service found that over 90% of participants spent less than 30 seconds on the privacy policy page (25; 164), and it seems that only 1% of people read the entire text of privacy policies (186).
Moreover, the lack of exercising rights is not limited to the right to information; it extends to other rights, including the right to erasure. Most Europeans (70% of survey respondents (144)) have never applied for data deletion, and only 1.13% frequently delete their cookie data (172). Service providers also report a lack of requests for data access and erasure (8; 231; 177). Some companies claim there had been no user access requests and less than ten profile deletion requests over a full year, contrary to their expectations (231; 177). Additionally, data portability is infrequently utilized; only 7% of users surveyed have exercised this right (186). Therefore, it is not common for users to exercise GDPR data rights.
@@
In contexts involving their children (e.g., smart homes), users can also pay more attention to privacy policies in smart devices surveilling their homes (223); certain parents consider privacy policies crucial to their children’s safety when selecting smart devices (223). Furthermore, some users regularly examine and delete personal data collected by smart devices related to their daily lives (130; 151; 44).
@@
Data rights have a significant influence on users’ perceived privacy concerns and uncertainties. Specifically, providing users with higher, simpler control over their data can mitigate their perceived privacy concerns, risks, and uncertainties (249; 95; 74; 128). A user experiment, for instance, demonstrates that showing users the personal information they have shared can decrease perceived privacy risks compared to merely viewing a privacy policy (95). When the system further allowed users to delete and modify their data, perceived privacy concerns and risks decreased.
@@
However, this observation contradicts previous findings (38; 156; 45) suggesting that providing users with data control can paradoxically lead to more data disclosure, a phenomenon known as the Peltzman effect where people tend to take riskier actions under security measures. While the exact reason for this contradiction is not clear, one hypothesis is that the mention of GDPR itself stimulated privacy-seeking behavior among users in the study (171). The only major difference across these studies is the mention of GDPR, and it’s possible that the priming effect of the term GDPR outweighed the Peltzman effect. Relatedly, one study (45) reports that granting data control to users did not significantly alter their data-sharing behavior; most participants maintained their willingness not to share personal information even after being given data control. Furthermore, evidence shows that users are more inclined to disclose their health data when given a reason for data collection that highlights the benefits of the product (24). In conclusion, further research is necessary to understand the impact of GDPR data rights on data disclosure behavior.
@@ 4. Companies’ perspective
A significant number of companies demonstrate a limited understanding of data rights (257; 159; 161). A user survey of businesses across eight EU countries revealed that only approximately 50-65% of participants correctly answered questions about the right to access, the right to erasure, and the right to object (257). Additionally, some app developers mistakenly believe that the GDPR is inapplicable to them because their primary market is outside the EU (159). Similarly, several startups misinterpret the GDPR, asserting that certain data rights cannot apply to their key technologies (161). Contrary to their beliefs, the GDPR applies to all businesses collecting personal data from citizens of European Economic Area (EEA) countries, irrespective of their primary market and business sectors. Notably, studies by Zanker et al. (257) and Nguyen et al. (159) conducted in early 2020 and early 2021, respectively, highlight the persistent deficiency in companies’ understanding of GDPR data rights even 2-3 years after its implementation.
@@
(C2.2) Willingness to implement data rights. Not all developers disregard users’ data rights. Many service providers emphasize the importance of enforcing them (100; 261; 9; 79; 143; 142; 86). They highlight the growing user awareness of their rights and believe the exercise of data rights will gain prominence with the adoption of GDPR (100; 261). Most European business leaders (about 86%) recognize the importance of GDPR compliance (79), a figure significantly higher than average individuals fully understanding GDPR data rights (50-65%) (257). This indicates that service providers can appreciate the importance of GDPR, even with a partial understanding of the law. Reasons for compliance include fear of fines and respect for users’ data rights (79). Likewise, practitioners knowledgeable in ML showed appreciation for the right to get a human review in automated decision-making and stated the importance of ML’s transparency (143; 142).
@@
Adherence to the GDPR data rights necessitates significant effort from companies, prompting various internal measures such as employee training, consultation procurement, and data management process modifications (257; 79; 97; 100). A survey reveals that approximately 91% of European companies have conducted at least one employee training session on this subject (257). Numerous companies report engaging in GDPR compliance consultation services (79). In preparation for the GDPR, most companies have incurred costs between €1,000 and €50,000, predominantly for employee education and consultancy services (257; 79).
@@
Several companies have implemented substantial changes to their data management processes to comply with data rights (97; 100; 111). An interview with a publicly listed ML company revealed significant changes to their data handling practices over six months, guided by their legal team to ensure GDPR compliance (97). Enterprise Architecture Management (EAM), which provides companies with strategies for efficient development, has played a crucial role in implementing data rights (100). For instance, some EAMs have defined procedures for developers to follow when implementing data rights and have specified the correct communication channels for notification.
@@
(C4.1) Right to information. A plethora of studies have scrutinized the compliance of privacy policies with the GDPR across various platforms such as web, mobile applications, and IoT (5; 93; 138; 256; 201; 237; 167; 220; 40; 73; 152; 231; 234; 155; 174; 224; 202; 10; 136; 260; 208; 225; 226; 165; 27; 6; 61; 26; 35; 98; 150; 122). These studies employ manual analyses or automated tools, such as Machine Learning (ML) or Natural Language Processing (NLP), to scrutinize privacy policies. Although methodologies and evaluation criteria vary across studies, the consensus is that few privacy policies currently meet all GDPR requirements.
@@
A noticeable disparity exists in the compliance rates for individual GDPR requirements. Over 90% of policies outline the collection of personal data and its purposes, though some instances of incomplete information have been reported (138; 256; 73; 260; 122). However, the majority of policies do not describe data rights like access and erasure (138; 35) even though European websites tend to state more extensive data rights than websites from other countries (252). Furthermore, many policies struggle to accurately inform users about the flow and transfer of data (234; 260; 221; 117; 150). A significant difference in compliance rates across websites in various countries is also found; European websites tend to have a higher compliance rate than others (252; 98).
@@
Surprisingly, some services do not provide a privacy policy at all (136; 208; 234; 166; 165; 150; 149). The availability of privacy policies varies significantly across platforms. For instance, most websites (4,566 out of 5,045 surveyed) provide a privacy policy (166). Yet, a mere 16% of 6,843 analyzed pornographic websites have a privacy policy (234). Similarly, only 62% of Google Assistant actions, which serve as voice-assistance applications, provide a privacy policy (136). Moreover, more than half of Greek and UK and more than one-third of German websites did not include a cookie notice (117; 125), and 12 out of 30 fertility apps analyzed did not prompt privacy notices (150).
@@
(C4.2) Rights to access, data portability, erasure, and object. A multitude of studies have delved into the extent to which systems uphold the right to access and data portability, typically by issuing a direct request (230; 238; 228; 42; 126; 216; 231; 232; 251; 262). These studies generally reveal that the majority of requests are honored within the one-month GDPR deadline. Similarly, an interview indicated that over 90% of service providers processed all requests for access and erasure within a month (177). However, compliance rates can be low in specific sectors (126; 232); for instance, a 2019 study showed that only 41% of popular German mobile apps provided a copy of personal data within a month (232).
@@
Moreover, past studies frequently observe that the data procured through the access right are incomplete (238; 228; 232; 126; 180); the data may not include certain personal information that users supplied. Furthermore, while many systems allow users to download their personal data in a machine-readable format, few provide the data import features required for data portability (127; 230). There are examples of systems rejecting users’ requests to download their data without providing a justification (262).
@@
Numerous technical and managerial challenges can make it virtually impossible for businesses to implement data rights. For instance, some suggest that, for efficiency, companies should deploy a system managing all applications processing personal information, as users can make daily requests to control their data (100). However, this would be challenging for applications run by independent departments (100). Furthermore, budget constraints can impede the implementation of data rights, as companies need to invest significantly in employee training and IT solutions for data management (65; 211).
(C5.1) Right to erasure. Among the eight GDPR rights, the right to erasure was most frequently cited as challenging by developers (161; 14; 144; 34; 97; 100; 47; 182). Removing user data, especially from backups and archives, appears to be a complex, costly, and unclear process (100; 47; 144). When backups need to be restored, previously deleted data may resurface, suggesting a failure in the erasure process. Moreover, larger enterprises tend to struggle more with this right due to the sheer volume of data (182). Some even admit that applying this right in their systems is impossible (14).
Confusion amplifies when considering ML (34; 34; 97). There is a lack of clear guidelines on whether to remove deleted data from all model training, test, and validation sets, and whether to delete the model itself upon receiving an erasure request from users. Another aspect is a conflict between data rights and technologies’ goals. Representatively, blockchain companies suffer from this problem (161). Given the immutability of blockchain technology, erasing data stored in the system is inherently impossible. Many blockchain companies discuss this issue in their privacy policies, with some stating outright that exercising this right is impossible (197). However, a few claims they can delete personal information from their blockchain systems, but details on how this is achieved are typically undisclosed (197).
@@
(C5.2) Right to information and access. Companies often find the rights to information and access challenging to implement (159; 14; 231). Firms must provide accurate information about data processing to users, necessitating a comprehensive understanding of data flows and processing mechanisms. However, many companies struggle with this, especially when third parties involved in online advertising also process the data. Full awareness of how these third parties handle user data can be a formidable task. Several studies reveal that most developers grapple with this lack of knowledge (159; 231; 133).
Moreover, the right to information poses a challenge in ML systems (97; 82; 62; 33). The right to information necessitates that corporations disclose the underlying logic of algorithms that may have consequential or legal effects on users. This disclosure, however, may not be feasible due to the inherent “black box” nature of ML algorithms. There are extant techniques, such as LIME (194), devised to demystify this issue, yet the resultant explanations are often technically complex and inaccessible to users lacking comprehensive ML knowledge (62; 33). This complexity often leads engineers to use these interpretability methods primarily to enhance ML system efficiency, deviating from the original intent of explicating algorithms to users (33).
Moreover, tensions may emerge between data rights and corporate interests, as ML algorithms can be considered proprietary to a company. This poses a critical question: To what extent should companies be obligated to reveal their algorithms (62)? A high degree of ML transparency could paradoxically infringe on user privacy by potentially revealing user identities embedded within the training dataset(62; 33).